OpenStudy (sleepyjess):

Analyze the function f(x) = - tan 4x. Include: - Domain and range - Period - Two Vertical Asymptotes

OpenStudy (blurbendy):

All I know is that the range is all real numbers. sorry, been a while since I've done this stuff.

ganeshie8 (ganeshie8):

\[\large \rm f(x) = -\tan(4x)\]

ganeshie8 (ganeshie8):

Say \(\large \rm 4x= t\) \[\large \rm f(t) = -\tan(t)\]

ganeshie8 (ganeshie8):

what do you know about domain of a normal `tan` function ?

OpenStudy (sleepyjess):

all real numbers except when cos x = 0

OpenStudy (sleepyjess):

@ganeshie8

ganeshie8 (ganeshie8):

Vrey good. if we write \(\large \tan (4x) = \dfrac{\sin (4x)}{\cos (4x)}\) can we say the domain of \(\tan(4x)\) is all real numbers except when \(\cos(4x) = 0\) ?

OpenStudy (sleepyjess):

how do i find the period?

ganeshie8 (ganeshie8):

whats the period of \(\tan (t)\) ?

OpenStudy (sleepyjess):

\(\pi/2\)

ganeshie8 (ganeshie8):

nope, try again

OpenStudy (sleepyjess):

\(\pi\)?

ganeshie8 (ganeshie8):

yes, since \(\tan (x)\) takes a time of \(\pi\) seconds to complete one cycle, \(\tan (4x)\) will take just \(\large \dfrac{\pi}{4}\) seconds to complete one cycle

ganeshie8 (ganeshie8):

so the period of tan(4x) would be pi/4

ganeshie8 (ganeshie8):

basically tan(4x) runs four times faster than tan(x), so its period(time) is also 4 times less

OpenStudy (sleepyjess):

would the asymptotes be \(\pi/8~and~-\pi/8\)?

ganeshie8 (ganeshie8):

Looks god!